# White Dwarf Structure and Binary Inspiral Gravitational Waves from Quantum Hadrodynamics

### v1

@@ Abstract
White dwarf, one of the compact objects in the universe, is not only pivotal for astrophysical inquiries but also furnishes a dynamic arena for nuclear physics exploration. In this work, we extend the relativistic mean field approach using a Walecka-type quantum hadrodynamics model to encapsulate the intricate structure of white dwarfs. This methodological advancement facilitates a cohesive analysis of white dwarfs, neutron stars, and the nuclear pasta within a unified theoretical framework. By meticulously calibrating the model parameters to nuclear matter properties, we successfully replicate the structures of various nuclei, such as ${}^{4}\rm He$, isotopes of $\rm C$, and ${}^{16}\rm O$. Subsequently, we predict the properties of white dwarfs composed of atom-like units and the gravitational waves stemming from binary white dwarf inspirals incorporating tidal deformability contributions up to the 2.5 post-Newtonian order. These results shed light on the structure of white dwarfs and offer more clues for future gravitational wave detection.
1 Introduction
@@
The detection of gravitational waves (GWs) by the LIGO/Virgo Collaborations Abbott et al. (2016, 2017, 2019) signifies that astronomy has entered the era of multi-messenger observations, providing a new means to probe the nature of compact objects such as black holes (BHs) and neutron stars (NSs). The space-based detectors like LISA Amaro-Seoane et al. (2017), Taiji Hu & Wu (2017), and TianQin Luo et al. (2016) which are scheduled to be launched in the 2030s aim to the GWs with frequency $0.1$mHz-$0.1$Hz. One of the sources of these space-based low-frequency GW detectors is the white dwarf (WD) binaries Ruiter et al. (2010); Yu & Jeffery (2010); Tamanini & Danielski (2019); McNeill et al. (2020); Wolz et al. (2020); Korol et al. (2020); Gibney (2024)—another kind of compact object in the universe.
@@
In addition to their astronomical significance, WDs have profound implications for nuclear physics. Binary WDs are intimately connected to Type Ia supernova (SN) explosions, events rich in physics involving strong interaction, weak interaction, and intense electromagnetic fields Niemeyer & Woosley (1997); Jose & Hernanz (1998); Fryer et al. (1999); Rosswog et al. (2009); Toonen et al. (2012); Katz & Dong (2012); Kashiyama et al. (2013). Understanding the structure of WDs, especially the components of the cores of WDs, is a complex yet crucial task for nuclear physics, and GWs from binary WD systems provide valuable insights into the characteristics of the internal structure of WDs and therefore the nuclear forces McNeill et al. (2020); Saumon et al. (2022).
@@
In the description of nuclear force, the quantum hadrodynamics (QHD) written in terms of colorless hadrons plays an indispensable role Walecka (1974); Serot & Walecka (1986); Weinberg (1990); Serot & Walecka (1997). In the realm of hadron interactions, linear and non-linear realizations of chiral symmetry stand out as frameworks that more directly mirror QCD properties Weinberg (1979, 1990); Gasser et al. (1988); Bando et al. (1985, 1988); Harada & Yamawaki (2003), offering an alternative to the traditional Walecka-type parametrization. With respect to the flavor or chiral symmetry, the electromagnetic and weak forces can be self-consistently included. Therefore, QHD is a unified framework for nucleon, nuclei, and their electromagnetic and weak processes.
@@
In this work, anchored on the Walecka-type QHD, we extend the relativistic mean field (RMF) approach which has been widely use in the study of homogeneous nuclear matter (NM) and NM clusters Watanabe et al. (2000); Horowitz et al. (2005); Xia et al. (2022); You et al. (2024) to study WD properties. The WDs are suggested to be made of charge-neutral objects—atom-like units—which have finite nucleon number nucleons within a box surrounded by an electron cloud. These objects effectively replicate the Wigner-Seitz cell structure of WDs Salpeter (1961); Rotondo et al. (2011a). The approach developed here can be taken as the microscopic materialization of the phenomenological parameterization of EOS of WD matter, such as the Thomas-Fermi model and Feynman-Metropolis-Teller (FMT) model Feynman et al. (1949); Rotondo et al. (2011b). The primary advantage of this extension is the ability to describe NSs, NS crusts, and WDs within a singular framework by utilizing a consistent set of interaction parametrizations, facilitating seamless transitions between modeling binary NSs, WDs, and NS-WD systems.
@@
Utilizing the above extension, we calculate the properties of $\rm{}^{4}He$, isotopes of $\rm C$ and $\rm{}^{16}O$ with the parameters of the model meticulously fitted by the properties of NM and the structural characteristics of NSs. Then, we deduce the structures of WDs, e.g., mass-radius (M-R) relation and tidal deformability (TD), composed of different elements ($\rm{}^{4}He$, $\rm{}^{12}C$, and $\rm{}^{16}O$). Finally, we study the GW signals stemming from the inspirals of the binary WDs with and without TD through the application of the post-Newtonian (PN) approximation Cutler et al. (1993); Damour et al. (2001) at the 2.5PN order. Our results of the WDs made of $\rm{}^{4}He$, isotopes of $\rm C$ and $\rm{}^{16}O$ are consistent with the observation data and Chandrasekhar limit, where free electron gas is assumed, and the GW signals stemming from the binary WDs composed of different elements are investigated. These results can serve as preliminary guides for future GW detection, offering insights into the nuances of binary WD dynamics.
@@
The rest of this paper is organized as follows: Sec. 2 introduces the framework of the extended RMF method and presents numerical results pertaining to the properties of NM and nuclei. Sec. 3 explores the structures of WDs and simulates the GW signals resulting from binary WD inspirals. Our discussion and outlook are given in the last section. We devote ourselves to the structure of nuclei and the EOS of WDs in Appendix A and the detail of PN expansion in Appendix B.
2 The extended RMF and nucleus
@@
Without loss of generality, we consider a Walecka-type model which has been widely used in nuclear physics Sugahara & Toki (1994); Shen et al. (1998). Including electromagnetic (EM) interaction, the model is written as:
@@
[…]
@@
${}-\frac{1}{4}\vec{\mathrm{P}}_{\mu\nu}\cdot\vec{\mathrm{P}}^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}\vec{\rho}_{\mu}\cdot\vec{\rho}^{\mu}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\ ,$
@@
with $\Psi=\left(\begin{array}[]{l}p\\ n\end{array}\right)$ and $\psi$ being, respectively, nucleon iso-doublet and electron fields. $A_{\mu}$ is the EM field, $\sigma,\omega_{\mu}$ and $\vec{\rho}_{\mu}=\rho_{\mu}^{i}\tau^{i}$ (with $\tau_{i}$ being the Pauli matrix) are isoscalar-scalar, isoscalar-vector and isovector-vector meson fields, respectively. The pseudoscalar mesons $\pi$ are neglected since they vanish in the RMF approximation. $\vec{\mathrm{P}}_{\mu\nu}=D_{\mu}\vec{\rho}_{\nu}-D_{\nu}\vec{\rho}_{\mu}$ is the field-strength tensor of rho meson fields. The covariant derivatives are defined as
$D_{\mu}\Psi$ $=$ $\left(\partial_{\mu}+iA_{\mu}Q\right)\Psi\ ,$
$D_{\mu}\psi$ $=$ $\left(\partial_{\mu}-ieA_{\mu}\right)\psi\ ,$
$D_{\mu}\vec{\rho}_{\nu}$ $=$ $\partial_{\mu}\vec{\rho}_{\nu}+iA_{\mu}\left[Q,\vec{\rho}_{\nu}\right]\ ,$ (3)
@@
By taking appropriate boundary conditions (BCs), the homogeneous NM and NM clusters can be calculated using the standard RMF approach  Watanabe et al. (2000); Horowitz et al. (2005); Xia et al. (2022). However, these BCs cannot be naively extended to the WD matter since the repulsive force from the EM interaction is much stronger than the attractive nuclear force. Here, we propose that the WD is made of atom-like units which have a nucleus core surrounded by homogeneous electron gas, as depicted in Fig. 1. For simplicity, we assume that both WD and its constitute unit are spherically symmetric.
Figure 1: The carton of WDs. WDs are composed of atom-like units with a single nuclei inside and the nuclei are surrounded by electron gas.
Then, homogeneous condition degenerates into isotopy so that the time derivatives and spatial components of meson fields can be neglected. The equations of motion (EOMs) to be solved are
@@
where $n_{i}$ and $n_{i}^{s}$ are, respectively, the number density and scalar density of particle “$i$", and the scalar density
@@
Note that in Eq. (2) only the zeroth component of the vector meson fields are considered and the component indices are omitted. By using the EOMs (2) and (2), the Hamiltonian (energy) density can be obtained via Legendre transformation from  (1) in a straightforward way.
To study the structure of nuclei inside the atom-like unit using model (1), we adopt the parameter values listed in Tab. 1 which were obtained by meticulously calibrating to align with the properties of NM and the M-R relations of NSs Guo et al. (2024). The nuclei are assumed to be spherical, and energy density can be calculated by ignoring the contribution from electrons. In our calculation, the initial distribution of fields are set to be within a sphere with radius $R=5\ \rm fm$, and the EOM (2) is iteratively solved to find the ground state of corresponding nuclei. Our results of the nuclei properties are shown in Tab. 2.
Table 1: Optimal values of the parameters obtained from pinning down NM properties and NS star structures Guo et al. (2024).
@@
$9.82$   $11.8$   $3.42$  $1.26$  $72.6$  $-1550$ MeV   $531$ MeV
{widetext}
Table 2: The results of nuclei properties. The root-mean-square radius $\sqrt{\langle r^{2}\rangle}$ is calculated via $\langle r^{2}\rangle=\int_{0}^{R_{0}}R^{2}\rho(R)4\pi R^{2}{\rm d}R/\int_{0}^{R_{0}}\rho(R)4\pi R^{2}{\rm d}R$, where $\rho(R)$ is the distribution of nucleons of the ground state and $R_{0}$ is the appropriate truncation for corresponding integral. The binding energy $\rm B.E.$ is defined by $m_{N}-E_{i}/({\rm N+Z})$, where “$i$" refers to different elements with $N(Z)$ being the number of neutron (proton) inside the nuclei. The empirical values (denoted as “Emp.") are taken from Refs. Fortune (2016); National Nuclear Data Center (2024); Pössel (2010). The radius $\sqrt{\langle r^{2}\rangle}$ and B.E. are in the units of fm and MeV, respectively.
Elements ${}^{4}\rm He$ ${}^{12}\rm C$ ${}^{13}\rm C$ ${}^{14}\rm C$ ${}^{15}\rm C$ ${}^{16}\rm C$ ${}^{17}\rm C$ ${}^{18}\rm C$; ${}^{19}\rm C$ ${}^{20}\rm C$ ${}^{16}\rm O$
$\sqrt{\langle r^{2}\rangle}$(Emp.) - - - $2.33(7)$ $2.54(4)$ $2.74(3)$ $2.76(3)$ $2.86(4)$ $3.16(7)$ $2.98(5)$
$\sqrt{\langle r^{2}\rangle}$(The.) $1.72$ $2.24$ $2.30$ $2.35$ $2.41$ $2.47$ $2.54$ $2.60$ $2.66$ $2.72$ $2.43$
B.E. (Emp.) $7.07$ $7.68$ $7.47$ $7.52$ $7.10$ $6.92$ $6.56$ $6.43$ $6.12$ $5.96$ $7.98$
B.E. (The.) $4.29$ $7.17$ $7.33$ $7.35$ $7.28$ $7.14$ $6.96$ $6.75$ $6.52$ $6.29$ $7.82$
From Tab. 2, one can see that our current approach successfully captures the qualitative properties of nuclei, except the B.E. of ${}^{4}\rm He$, which is unnaturally deep bounded, see Ref. Beane et al. (2013). Additionally, the results increasingly diverge from empirical data with $N-Z$. This deviation is attributed to the oversimplified parametrization of isospin-related interactions, especially those involving the $\rho$ meson.
3 The structure of white dwarfs and binary inspirals
The energy density $\mathcal{E}$ of the atom-like units can be calculated by using model (1) and pressure can be obtained through

### v2

@@ Abstract
White dwarfs, one of the compact objects in the universe, play a crucial role in astrophysical research and provide a platform for exploring nuclear physics. In this work, we extend the relativistic mean field approach by using a Walecka-type quantum hadrodynamics model to capture the intricate structure of white dwarfs. We calculate nuclear properties, Coulomb energy, and photon energy within white dwarfs in a unified framework. By carefully calibrating the model parameters to align with nuclear matter properties, we successfully reproduce the structures of several elements in white dwarfs, such as the isotopes of $\rm C$ and $^{16}\rm O$, except for the unnaturally deeply bound state 4He. Furthermore, we predict the characteristics of white dwarfs composed of atom-like units and the gravitational waves stemming from binary white dwarf inspirals incorporating tidal deformability contributions up to the 2.5 post-Newtonian order. These results shed light on the structure of white dwarfs and provide valuable information for future gravitational wave detection. This methodological advancement allows for a cohesive analysis of white dwarfs, neutron stars, and the nuclear pasta within a unified theoretical framework.
I Introduction
@@
The detection of gravitational waves (GWs) by the LIGO/Virgo Collaborations [1, 2, 3] signifies that astronomy has entered the era of multi-messenger observations. This advancement provids a new means to investigate the nature of compact objects such as black holes (BHs) and neutron stars (NSs). The space-based detectors like LISA [5], Taiji [26], and Tianqin [35] which are scheduled to be launched in the 2030s aim to the GWs with frequency 0.10.1mHz-0.10.1Hz. One of the sources of these space-based low-frequency GW detectors is the white dwarf (WD) binaries [50, 70, 40, 66, 31, 21]—another kind of compact object in the universe.
As one of the compact objects in the universe, WDs are stabilized by the equilibrium between the pressure of the degenerate electron gas and inner gravity. Due to the charge neutrality, in addition to electrons, there should be nuclei in WDs. In the pioneer work by Chandrasekhar [13], the WD matter was assumed to be the uniformed degenerate electron gas and pointlike nuclei. Then, by introducing the concept of Wigner-Seitz cell, Salpeter studied the corrections due to the non-uniformity of the electrons inside a Wigner-Seitz cell and found that the Coulomb force is essential for determining the maximum stable mass of non-rotating WDs [51]. In Refs. [48, 49], a relativistic Feynman-Metropolis-Teller model was considered to include the finite size of nuclei by assuming a constant distribution of protons confined in a radius characterized by the pion Compton wavelength. This model allows for a more accurate calculation of the energy and pressure of the Wigner-Seitz and consequently a more accurate equation of state (EOS) of WD matter. In the WD matter, the pressure is dominated by the degenerate electron gas, while the energy of a Wigner-Seitz cell is mainly from nucleus, and its binding energy is $10\rm MeV$ which is already larger than the contribution from electromagnetical dynamics. Therefore, a realistic description of the nucleus, radius and binding energy, requiring an accurate understanding of the nuclear force which is crucial for the study of WDs.
@@
Binary WDs are closely connected to Type Ia supernova (SN) explosions, which are events rich in physics, involving strong interaction, weak interaction, and intense electromagnetic fields [43, 19, 47, 61]. Understanding the structure of WDs, especially the components of the cores of WDs, is a complex yet crucial task for nuclear physics. GWs stemming from binary WD systems provide valuable insights into the characteristics of the internal structure of WDs and therefore the nuclear forces [40, 52].
@@
In the description of nuclear force, quantum hadrodynamics (QHD) written in terms of colorless hadrons plays an indispensable role [63, 54, 64, 55]. In the realm of hadron interactions, linear and non-linear realizations of chiral symmetry both are frameworks that more directly mirror QCD properties [64, 65, 20], providing an alternative to the traditional Walecka-type parametrization. With respect to the flavor or chiral symmetry, the electromagnetic and weak forces can be self-consistently included. Therefore, QHD serves as a unified framework for nucleons, nuclei, and their electromagnetic and weak processes.
@@
In this work, anchored on the Walecka-type QHD including hadrons, electrons and photons, we extend the relativistic mean field (RMF) approach which has been widely used in the study of homogeneous nuclear matter (NM) and NM clusters [68, 67] to study WD properties. The WDs are suggested to be made of charge-neutral objects—atom-like units—which have finite nucleon number nuclei surrounded by an electron cloud. These objects effectively replicate the Wigner-Seitz cell structure of WDs [51, 49] but with the extension of the nucleus calculated from the Walecka-type model. The approach developed here can be taken as a microscopic materialization of the phenomenological parametrization of EOS of WD matter and a unified framework for NSs, NS crusts, and WDs. Note that we will not consider the strange degrees of freedom which are found to shrink the size of white dwarfs [32] since it is beyond the purpose this work.
@@
Using the aforementioned extension, we calculate the properties of $\rm{}^{4}He$, isotopes of $\rm C$ and $\rm{}^{16}O$ with model parameters carefully fitted by the properties of NM and the structures of NSs. We then derive the structures of WDs, including their mass-radius (M-R) relations and tidal deformability (TD), for various compositions ($\rm{}^{4}He$, $\rm{}^{12}C$, and $\rm{}^{16}O$). Finally, we study the GW signals stemming from the inspirals of the binary WDs with and without TD through the application of the post-Newtonian (PN) approximation [14, 15] at the 2.5PN order. Our results of WDs made of $\rm{}^{4}He$, isotopes of $\rm C$ and $\rm{}^{16}O$ are consistent with the observations and Chandrasekhar limit [13], which assumes a free electron gas. We also investigate the GW signals stemming from the binary WDs composed of different elements. These results may serve as preliminary guides for future GW detection, providing insights into the intricacies of binary WD dynamics.
@@
The rest of this paper is organized as follows: Sec. II introduces the framework of the extended RMF method and presents numerical results pertaining to the properties of NM and nuclei. Sec. III explores the structures of WDs and simulates the GW signals resulting from binary WD inspirals. Our discussion and outlook are given in the last section. We describe the structures of nuclei and the EOS of WDs in Appendix A and the detail of PN expansion in Appendix B.
II The extended RMF and nucleus
@@
Without loss of generality, we consider a Walecka-type model which has been widely used in nuclear physics [58, 56]. Including electromagnetic (EM) interaction, the model is written as:
@@
$\mathcal{L}_{\rm fermion}$ $=$ $\bar{\Psi}\left(i\not{\text{D}}-m_{N}\right)\Psi+\bar{\psi}\left(i\not{\text{D}}-m_{e}\right)\psi\ ,$
@@
${}-\frac{1}{4}\vec{\bm{P}}_{\mu\nu}\cdot\vec{\bm{P}}^{\mu\nu}+\frac{1}{2}m_{\rho}^{2}\vec{\rho}_{\mu}\cdot\vec{\rho}^{\mu}-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}\ ,$
@@
with $\Psi=\left(\begin{array}[]{l}p\\ n\end{array}\right)$ and $\psi$ being, respectively, nucleon iso-doublet and electron fields. $A_{\mu}$ is the EM field, $\sigma,\omega_{\mu}$ and $\vec{\rho}_{\mu}=\rho_{\mu}^{i}\tau^{i}$ (with $\tau_{i}$ being the Pauli matrix) are isoscalar-scalar, isoscalar-vector and isovector-vector meson fields, respectively. The pseudoscalar mesons $\pi$ are neglected since they vanish in the RMF approximation. $\vec{\bm{P}}_{\mu\nu}=\text{D}_{\mu}\vec{\rho}_{\nu}-\text{D}_{\nu}\vec{\rho}_{\mu}-2ig_{\rho}\vec{\rho}_{\mu}\times\vec{\rho}_{\nu}$ is the field-strength tensor of rho meson fields. The covariant derivatives are defined as
$\text{D}_{\mu}\Psi$ $=$ $\left(\partial_{\mu}+iA_{\mu}Q\right)\Psi\ ,$
$\text{D}_{\mu}\psi$ $=$ $\left(\partial_{\mu}-ieA_{\mu}\right)\psi\ ,$
$\text{D}_{\mu}\vec{\rho}_{\nu}$ $=$ $\partial_{\mu}\vec{\rho}_{\nu}+iA_{\mu}\left[Q,\vec{\rho}_{\nu}\right]\ ,$ (3)
@@
By taking appropriate boundary conditions (BCs), the homogeneous NM and NM clusters can be calculated using the standard RMF approach [25, 67]. However, these BCs cannot be naively extended to the WD matter, as the repulsive force from the EM interaction is much stronger than the attractive nuclear force. Here, we propose that the WD is universally made of atom-like units—Wigner Seitz cells—each containing a nucleus at its core, surrounded by homogeneous electron gas, as illustrated in Fig. 1. The nucleus is calculated by using the microscopic Walecka-type model (1) therefore has a realistic internal structure. The electromagnetic force is obtained by solving the Maxwell equation with a proton source. For simplicity, we assume that both WD and its constituent units are spherically symmetric.
Figure 1: The carton of WDs. WDs are composed of atom-like units with a single nucleus inside and the nucleus is surrounded by electron gas.
Under the above conditions, the homogeneous condition simplifies to isotropy, allowing us to neglect the time derivatives and spatial components of the meson fields. Furthermore, after applying the RMF approximation, only the neutral $\rho$ meson survives. Consequently, the equations of motion (EOMs) that need to be solved are:
@@
where $n_{i}$ and $n_{i}^{s}$ are, respectively, the number density and scalar density of particle “ii", and the scalar density
@@
Note that in Eq. (II) only the zeroth component of the vector meson fields are considered, and the component indices are omitted. By using the EOMs (II) and (II), the Hamiltonian (energy) density can be obtained via a straightforward Legendre transformation from (1). It can be decomposed as
$\mathcal{E}$ $=$ $\mathcal{E}_{\rm N}+\mathcal{E}_{e}+\mathcal{E}_{\rm NM}+\mathcal{E}_{\rm p\gamma}+\mathcal{E}_{\rm e\gamma}+\mathcal{E}_{\gamma}+\mathcal{E}_{M},$ (7)
where $\mathcal{E}_{\rm N},\ \mathcal{E}_{\rm e},\ \mathcal{E}_{\rm NM},\ \mathcal{E}_{\rm p\gamma},\ \mathcal{E}_{\rm e\gamma},\ \mathcal{E}_{\gamma}$ and $\mathcal{E}_{\rm M}$ represent, respectively, the contributions from free nucleon with medium modified mass $m_{N}^{\ast}$, free electron, nucleon-meson interaction, proton-photon interaction, electron-photon interaction, pure photon kinetic energy and pure meson interaction. Explicitly,
$\mathcal{E}_{\rm N}$ $=$ $\sum_{i=n,p}\frac{1}{8}\left[t_{i}\sqrt{1+t_{i}^{2}}\left(1+2t_{i}^{2}\right)-\operatorname{arcsinh}t_{i}\right],$
$\mathcal{E}_{\rm e}$ $=$ $\frac{1}{8}\left[t_{e}\sqrt{1+t_{e}^{2}}\left(1+2t_{e}^{2}\right)-\operatorname{arcsinh}t_{e}\right]\ ,$
$\mathcal{E}_{\rm NM}$ $=$ $\Psi^{\dagger}\left(g_{\omega}\omega+g_{\rho}\rho\right)\Psi\ ,$
$\mathcal{E}_{\rm p\gamma}$ $=$ $A\Psi^{\dagger}Q\Psi\ ,$
$\mathcal{E}_{\rm e\gamma}$ $=$ $-\psi^{\dagger}eA\psi\ ,$
$\mathcal{E}_{\gamma}$ $=$ ${}-\frac{1}{2}\left(\nabla A\right)^{2}\ ,$
$\mathcal{E}_{\rm M}$ $=$ ${}-\frac{1}{2}\left(\nabla\sigma\right)^{2}-\frac{1}{2}m_{\sigma}^{2}\sigma^{2}+\frac{1}{3}g_{2}\sigma^{3}+\frac{1}{4}g_{3}\sigma^{4}$ (8)
${}-\frac{1}{2}\left(\nabla\omega\right)^{2}-\frac{1}{2}m_{\omega}^{2}\omega^{2}+\frac{1}{4}c_{3}\omega^{4}$
${}-\frac{1}{2}\left(\nabla\rho\right)^{2}-\frac{1}{2}m_{\rho}^{2}\rho^{2}\ .$
To study the structure of nuclei within the atom-like units using model (1), we adopt the parameter values listed in Tab. 1 which were obtained by meticulously calibrating to align with the properties of NM and the M-R relations of NSs [23]. The nuclei are assumed to be spherical, and the energy density can be calculated by neglecting the contributions from electrons within each atom-like unit, as a result of Gauss theorem, while still including the the EM interactions between protons. Specifically, contributions $\mathcal{E}_{\rm N},\mathcal{E}_{\rm NM},\mathcal{E}_{\rm p\gamma},\mathcal{E}_{\gamma}$ and $\mathcal{E}_{M}$ from Eq. (7) are considered. In our calculations, the initial distribution of hadron fields is set within a sphere of radius $R_{0}=5\ \rm fm$, which is a typical size of a nucleus. The EOMs (II) are iteratively solved to determine the ground state of corresponding nuclei, which gives us the reasonable geometry information of the corresponding nuclei. The boundary conditions are defined as the nuclear density $n_{n(p)}(r=R_{0})=0$ and $n^{\prime}_{n(p)}(r=R_{0})=0$. Explicit check using $R_{0}=(4-10)$ fm shows that the results of nucleus properties are intact. Our results of the nuclei properties are shown in Tab. 2.
Table 1: Optimal values of the parameters obtained from pinning down NM properties and NS star structures [23].
@@
$9.82$   $11.8$   $3.42$  $1.26$  $72.6$  1550-1550 MeV   $531$ MeV
Table 2: The results of nuclei properties. The root-mean-square radius $\sqrt{\langle r^{2}\rangle}$ is calculated via $\langle r^{2}\rangle=\int_{0}^{R_{0}}R^{2}\rho(R)4\pi R^{2}{\rm d}R/\int_{0}^{R_{0}}\rho(R)4\pi R^{2}{\rm d}R$, where $\rho(R)$ is the distribution of nucleons of the ground state. The binding energy $\rm B.E.$ is defined by $m_{N}-E_{i}/(N+Z)$, where $E_{i}$ is the total energy of an atomic unit and “ii" refers to different elements with $N(Z)$ being the number of neutron (proton) inside the nuclei. The experiment values (denoted as “Exp.") are taken from Refs. [6, 18, 42]. The radius $\sqrt{\langle r^{2}\rangle}$ and B.E. are in units of fm and MeV, respectively.
Elements $^{4}\rm He$ $^{12}\rm C$ $^{13}\rm C$ $^{14}\rm C$ $^{15}\rm C$ $^{16}\rm C$ $^{17}\rm C$ $^{18}\rm C$ $^{19}\rm C$ $^{20}\rm C$ $^{16}\rm O$
$\sqrt{\langle r^{2}\rangle}$(Exp.) - - - $2.33(7)$ $2.54(4)$ $2.74(3)$ $2.76(3)$ $2.86(4)$ $3.16(7)$ $2.98(5)$ -
$\sqrt{\langle r^{2}\rangle}$(The.) $1.90$ $2.34$ $2.36$ $2.43$ $2.48$ $2.55$ $2.61$ $2.66$ $2.71$ $2.77$ $2.51$
B.E. (Exp.) $7.07$ $7.68$ $7.47$ $7.52$ $7.10$ $6.92$ $6.56$ $6.43$ $6.12$ $5.96$ $7.98$
B.E. (The.) $3.45$ $6.43$ $6.63$ $6.65$ $6.61$ $6.50$ $6.37$ $6.20$ $6.01$ $5.81$ $7.08$
From Tab. 2, one can see that our current approach successfully captures the qualitative properties of nuclei, with the exception of the binding energy of $^{4}\rm He$, which is unnaturally deep bounded [7]. Furthermore, the results increasingly diverge from experiment data with NZN-Z. This deviation is attributed to the oversimplified parametrization of isospin-related interactions, especially those involving the $\rho$ meson.
III Structure of white dwarfs and binary inspirals
The energy of an atom-like unit consists of the contributions from the nucleus core, which is the nucleus mass calculated above, as well as the energy of free electrons, free photons, and the Coulomb energy between the electrons and the nucleus core. By varying the volume of the unit, we can obtain both the the number density $\mathcal{N}$ and the energy density $\mathcal{E}$ of the atom-like units. The pressure of the unit can be derived through the thermal relation